# Towards Robust ESG Analysis Against Greenwashing Risks: Aspect-Action Analysis with Cross-Category Generalization

### v1

@@ Abstract
Sustainability reports are key for evaluating companies’ environmental, social and governance (ESG) performance, but their content is increasingly obscured by greenwashing - sustainability claims that are misleading, exaggerated, and fabricated. Yet, existing NLP approaches for ESG analysis lack robustness against greenwashing risks, often extracting insights that reflect misleading or exaggerated sustainability claims rather than objective ESG performance. To bridge this gap, we introduce A3CG - Aspect-Action Analysis with Cross-Category Generalization, as a novel dataset to improve the robustness of ESG analysis amid the prevalence of greenwashing. By explicitly linking sustainability aspects with their associated actions, A3CG facilitates a more fine-grained and transparent evaluation of sustainability claims, ensuring that insights are grounded in verifiable actions rather than vague or misleading rhetoric. Additionally, A3CG emphasizes cross-category generalization. This ensures robust model performance in aspect-action analysis even when companies change their reports to selectively favor certain sustainability areas. Through experiments on A3CG, we analyze state-of-the-art supervised models and LLMs, uncovering their limitations and outlining key directions for future research.
@@ 1 Introduction
Sustainability reports have become an important mechanism for evaluating a company’s environmental, social, and governance (ESG) performance Nguyen (2020). However, the rise of greenwashing - the practice of providing misleading, exaggerated, and fabricated sustainability claims - has undermined the credibility of sustainability reports Ong et al. (2024). By obscuring the evaluation of a company’s ESG performance, greenwashing hinders meaningful progress toward sustainability goals Rajesh (2020).
@@
Figure 1: Overview of the A3CG task
To tackle this, we propose A3CG - Aspect-Action Analysis with Cross-Category Generalization, a novel dataset for robust ESG analysis amid the prevalence of greenwashing. A3CG provides a foundation for systematically extracting two interconnected components from sustainability texts: i) Aspect: The sustainability-related entity, goal, or activity discussed (i.e. “carbon emissions reduction”). ii) Action: The type of engagement related to the aspect, “implemented”, “planning”, or “indeterminate”. While “implemented” and “planning” indicate concrete steps taken or planned toward the sustainability aspect, “indeterminate” highlights vague, evasive or non-attributable claims which may be linked with greenwashing Siano et al. (2017). By enabling the identification of sustainability aspects and their action assessments, A3CG allows models to clarify genuine sustainability initiatives from ambiguous or misleading rhetoric. While A3CG is not designed for definitive greenwashing detection, it facilitates clear and transparent ESG analysis that is less susceptible to greenwashing risks.
@@ 3 Dataset Construction
This section describes A3CG’s [^1] For full reproducibility, the dataset and all experimental codes will be released after the review process. Company names and the corresponding years of sustainability reports will also be available. Using this information, these public reports can be searched and accessed online. construction.
@@ 3.1 Data Collection & Quality Control
We collect sustainability statements from 1679 sustainability reportsfootnotemark: of Singapore Exchange (SGX) companies, from 2017-2022 inclusive. The reports are public and can be found online via company websites. For quality control, we eliminate: (1) Incomplete statements (2) Incoherent statements with misspelled words (3) Non-English statements.
@@ 3.2 Annotation Scheme
Pairs of (aspect, action) are annotated for each sustainability statement. In a pair, the action characterizes the company’s engagement with the aspect. Notably, the aspect’s sustainability category is also annotated. However, similar to other generalization studies Xu et al. (2023a), the aspect category is not a prediction target. Instead, it is utilized to split the dataset for cross-category testing. We summarize the aspect and action definitions [^2] Detailed annotation instructions, guidelines, definitions, samples, and annotator details can be found in appendix A below:
@@ 4.2 Models
SOTA-ABSAfootnotemark: : We adapt SOTA-ABSA models for A3CG. InstructABSA Scaria et al. (2023), IT-RER-ABSA Zheng et al. (2024), GRACE Luo et al. (2020), CONTRASTE Mukherjee et al. (2023).
Large Language Modelsfootnotemark: : The latest LLMs (zero-shot and few-shot) are evaluated. GPT-4o OpenAI (2024), Claude 3.5 Sonnet Huang et al. (2024b), Llama 3 (70B) AI@Meta (2024), DeepSeek V3 Liu et al. (2024).
@@ 5 Results & Discussion
From Table 2, model evaluation on the full dataset indicates the dataset’s stability, with supervised models (T5, T5+CL, CONTRASTE) achieving over 70% F1. In the following, we focus on A3CG’s core objective: aspect-action analysis (AAA) - the extraction of aspect-action pairs in a cross-category generalization setup. Therefore, we primarily discuss performance on unseen (US) categories. Performance on sub-tasks of AAA is also analyzed: aspect term extraction (ATE) for identifying aspects, and action classification (AC) for classifying action labels after correctly identifying aspects (AC excludes no-aspect cases).
5.1 Comparison of Model Types
While supervised models have higher absolute performance than LLMs, they show significant inefficiency at transferring performance from seen to unseen categories. For supervised models, the difference between average F1 scores on seen and unseen categories, $\Delta$, is significant - greater than 14% in magnitude except BERT-ST+AL ($\Delta$=-6.95%), from Table 2. This represents a substantial drop in performance when transferring learned patterns from seen to unseen categories, underscoring the weak generalization efficiency of learned features. Consequently, supervised models must improve the generalizability efficiency of their learned features, to enhance cross-category generalization.
Specific LLMs can still perform competitively on unseen categories. Despite being generally outperformed by supervised models, specific LLMs achieve competitive average F1 scores on unseen categories, narrowing the performance gap. From Figure 3, Claude 3.5 Sonnet+FS (42.03) outperforms several supervised models - T5 (41.12), BERT-ST (26.06), BERT-ST+CL (37.93), BERT-ST+AL (26.24), while lagging behind IT-RER-ABSA (43.26), T5+AL (42.82) by less than 1.5%. Therefore, while supervised learning provides optimal performance, its superiority is not absolute - specific LLMs can still present viable alternatives to full-scale supervised fine-tuning, particularly in low-resource scenarios.
Action classification is more challenging than aspect term extraction for all methods. From Figure 3, AC F1 scores are markedly lower than ATE on unseen categories across all methods. This is potentially because while both subtasks involve semantic, syntactic and pragmatic reasoning, AC places greater emphasis on interpreting complex syntax and pragmatics - an issue explored later. Hence, AC is particularly challenging for supervised and LLM approaches. Therefore, given the considerable difficulty and lower performance, targeted model improvements for AC could drive significant gains in the overall performance of AAA.
@@ 6 Conclusion
In this study, we proposed the A3CG dataset, and evaluated SOTA supervised models and LLMs to highlight key directions for research. We invite researchers to unlock AI’s potential for robust ESG analysis, by developing models tailored to A3CG.
@@
The A3CG dataset focuses on sustainability statements and reports that are solely in the English language. However, we acknowledge that corporate sustainability is a global responsibility that can encompass many different countries and regions. Therefore, for greater inclusivity, future work will focus on extending A3CG to non-English languages.

### v2

@@ Abstract
Sustainability reports are key for evaluating companies’ environmental, social and governance (ESG) performance. To analyze these reports, NLP approaches can efficiently extract ESG insights at scale. However, even the most advanced NLP methods lack robustness against ESG content that is greenwashed – i.e. sustainability claims that are misleading, exaggerated, and fabricated. Accordingly, existing NLP approaches often extract insights that reflect misleading or exaggerated sustainability claims rather than objective ESG performance. To tackle this issue, we introduce A3CG - Aspect-Action Analysis with Cross-Category Generalization, as a novel dataset to improve the robustness of ESG analysis amid the prevalence of greenwashing. By explicitly linking sustainability aspects with their associated actions, A3CG facilitates a more fine-grained and transparent evaluation of sustainability claims, ensuring that insights are grounded in verifiable actions rather than vague or misleading rhetoric. Additionally, A3CG emphasizes cross-category generalization. This ensures robust model performance in aspect-action analysis even when companies change their reports to selectively favor certain sustainability areas. Through experiments on A3CG, we analyze state-of-the-art supervised models and LLMs, uncovering their limitations and outlining key directions for future research.
@@ 1 Introduction
Sustainability reports have become an important mechanism for evaluating a company’s environmental, social, and governance (ESG) performance Nguyen (2020). To facilitate the analysis of these reports, NLP methods typically offer an automated and efficient way to extract ESG insights at scale Ong et al. (2024). However, the rise of greenwashing – the practice of providing misleading, exaggerated, and fabricated sustainability claims – has increasingly obscured the content of these reports Ong et al. (2024). Greenwashing presents a major challenge for NLP-driven ESG analysis, let alone manual assessments, ultimately hindering meaningful progress toward sustainability goals Rajesh (2020).
@@
Figure 1: A3CG Task Overview. Given a sustainability statement, the model extracts aspect–action pairs. To evaluate cross-category generalization, it is tested on statements containing aspect categories not seen during training – i.e. trained on "source", tested on "target".
To tackle this, we propose A3CG [^1] We release the dataset and experimental codes at:
https://github.com/keanepotato/a3cg_greenwash
- Aspect-Action Analysis with Cross-Category Generalization, a novel dataset for robust ESG analysis amid the prevalence of greenwashing. A3CG provides a foundation for systematically extracting two interconnected components from sustainability texts: i) Aspect: The sustainability-related entity, goal, or activity discussed (i.e. “carbon emissions reduction”). ii) Action: The type of engagement related to the aspect, “implemented”, “planning”, or “indeterminate”. While “implemented” and “planning” indicate concrete steps taken or planned toward the sustainability aspect, “indeterminate” highlights vague, evasive or non-attributable claims which may be linked with greenwashing Siano et al. (2017). By enabling the identification of sustainability aspects and their action assessments, A3CG allows models to clarify genuine sustainability initiatives from ambiguous or misleading rhetoric. Through this flagging of clear sustainability initiatives, stakeholders can then verify and track them more easily. Therefore, while A3CG is not designed for definitive greenwashing detection, it facilitates clear and transparent ESG analysis that is less susceptible to greenwashing risks.
@@ 3 Dataset Construction
This section describes A3CG’s construction.
@@ 3.1 Data Collection & Quality Control
We collect sustainability statements from 1679 sustainability reports [^2] Company names and the corresponding years of sustainability reports are also available at: https://github.com/keanepotato/a3cg_greenwash. Using this information, these public reports can be searched and accessed online. of Singapore Exchange (SGX) companies, from 2017-2022 inclusive. The reports are public and can be found online via company websites. For quality control, we eliminate: (1) Incomplete statements (2) Incoherent statements with misspelled words (3) Non-English statements.
@@ 3.2 Annotation Scheme
Pairs of (aspect, action) are annotated for each sustainability statement. In a pair, the action characterizes the company’s engagement with the aspect. Notably, the aspect’s sustainability category (i.e. "emissions control" as a category of the aspect "carbon emissions") is also annotated. However, similar to other generalization studies Xu et al. (2023a), the aspect category is not a prediction target. Instead, it is utilized to split the dataset for cross-category testing. We summarize the aspect and action definitions [^3] Detailed annotation instructions, guidelines, definitions, samples, and annotator details can be found in appendix A below:
@@ 4.2 Models
SOTA-ABSAfootnotemark: : We adapt SOTA-ABSA models for A3CG, given that both ABSA and A3CG tasks overlap in their focus on aspect-level analysis. Therefore, we hypothesize that ABSA methods may be applicable to A3CG, and test this through experiments. We select the latest sequence-tagging methods–GRACE Luo et al. (2020), and generative methods–InstructABSA Scaria et al. (2023), IT-RER-ABSA Zheng et al. (2024), CONTRASTE Mukherjee et al. (2023).
Large Language Modelsfootnotemark: : The latest LLMs (zero-shot and few-shot) are evaluated. These are selected from different LLM families, and have achieved state of the art performance on general reasoning and natural language understanding benchmarks. The LLMs evaluated include–GPT-4o OpenAI (2024), Claude 3.5 Sonnet Huang et al. (2024b), Llama 3 (70B) AI@Meta (2024), DeepSeek V3 Liu et al. (2024).
@@ 5 Results & Discussion
From Table 2, model evaluation on the full dataset indicates the dataset’s stability, with supervised models (T5, T5+CL, CONTRASTE) achieving over 70% F1. In the following, we focus on A3CG’s core objective: aspect-action analysis (AAA) - the extraction of aspect-action pairs in a cross-category generalization setup. Therefore, we primarily discuss performance on unseen (US) categories across all folds. Performance on sub-tasks of AAA is also analyzed: aspect term extraction (ATE) for identifying aspects, and action classification (AC) for classifying action labels after correctly identifying aspects (AC excludes no-aspect cases).
For the models evaluated in our paper, we distinguish between two types of models: supervised models and large language models (LLMs). Supervised models include, from Section 4.2–vanilla models, vanilla + learning paradigm and SOTA-ABSA. These models are based on pre-trained language models (PLMs) (i.e. T5, BERT), which are typically smaller in scale and fine-tuned for specific tasks (in our case, the A3CG task). In contrast, when we refer to LLMs, we refer to larger, general-purpose models (i.e. GPT-4o) that are trained on broad language tasks, and are implemented here via prompt-based methods without any fine-tuning. In the following, we discuss the results of our experiments, by first analyzing the performance of supervised methods (Sections 5.15.2), followed by LLMs (Sections 5.35.4). At the end, we compare both supervised methods and LLMs (Section 5.5) to contrast the strengths and weaknesses of both methods for A3CG.
@@ 5.4 A3CG Challenges for LLMs
5.5 Comparison of Model Types
While supervised models have higher absolute performance than LLMs, they show significant inefficiency at transferring performance from seen to unseen categories. For supervised models, although they demonstrate the best absolute performance on unseen categories, the difference between average F1 scores on seen and unseen categories, $\Delta$, is significant - greater than 14% in magnitude except BERT-ST+AL ($\Delta$=-6.95%), from Table 2. This represents a substantial drop in performance when transferring learned patterns from seen to unseen categories, underscoring the weak generalization efficiency of learned features. The features learned by LLMs are shaped by broad pre-training, making them generally applicable across diverse data distributions. As a result, the performance drop on unseen categories is relatively modest, as these categories do not represent truly out-of-distribution data. In contrast, supervised models rely on features tuned to the training distribution, and their performance shortfall on unseen categories reflects a stronger sensitivity to distribution shifts. Consequently, supervised models can improve the generalization efficiency of their learned features, to enhance cross-category generalization.
Specific LLMs can still perform competitively on unseen categories. Despite being generally outperformed by supervised models, specific LLMs achieve competitive average F1 scores on unseen categories, narrowing the performance gap. From Figure 3, Claude 3.5 Sonnet+FS (42.03) outperforms several supervised models - T5 (41.12), BERT-ST (26.06), BERT-ST+CL (37.93), BERT-ST+AL (26.24), while lagging behind IT-RER-ABSA (43.26), T5+AL (42.82) by less than 1.5%. Therefore, while supervised learning provides optimal performance, its superiority is not absolute - specific LLMs can still present viable alternatives to full-scale supervised fine-tuning, particularly in scenarios where computational resources are limited but API-based access to LLMs are available.
Action classification is more challenging than aspect term extraction for all methods. From Figure 3, AC F1 scores are markedly lower than ATE on unseen categories across all methods. This is potentially because while both subtasks involve semantic, syntactic and pragmatic reasoning, AC places greater emphasis on interpreting complex syntax and pragmatics, as discussed in the failure cases in supervised models (Section 5.1) and LLMs (Section 5.4). Hence, AC is particularly challenging for supervised and LLM approaches. Therefore, given the considerable difficulty and lower performance, targeted model improvements for AC could drive significant gains in the overall performance of AAA.
@@ 6 Conclusion
In this study, we discussed the greenwashing phenomenon, and highlighted significant limitations of current NLP methods for robust ESG analysis. To address this, we proposed the A3CG dataset and evaluated supervised models, and LLMs under zero and few-shot settings. While supervised models achieved the highest absolute performance for unseen category generalization, LLMs exhibited greater generalization efficiency from seen to unseen categories. Within supervised approaches, contrastive learning proved more effective than adversarial learning for unseen category generalization. Yet both supervised and LLM approaches face limitations. On the one hand, supervised models lack contextual awareness for detecting aspects of unseen categories and struggle to classify actions in complex syntactic cases. On the other hand, LLMs overlook non-environmental aspects possibly due to pretraining biases, and struggle with classifying actions where there is pragmatic nuance. These results underscore the need for further research into models that can generalize robustly across ESG categories for the aspect-action analysis task. Therefore, we invite researchers to experiment with A3CG to unlock the potential of NLP for robust ESG analysis.
@@
The A3CG dataset focuses on sustainability statements that are solely in English. However, we acknowledge that sustainability is a global responsibility that can encompass many different regions. Therefore, for greater inclusivity, future work will extend A3CG to non-English languages. Additionally, due to computational resources, this paper does not consider supervised fine-tuning of smaller LLMs–i.e. Llama 3 8B AI@Meta (2024). Yet, LoRA-based fine-tuning of these smaller scale LLMs has proven effective in other domain-specific tasks Varshney et al. (2025), and could represent a future research direction.
@@
Acknowledgements
This research/project is supported by the NUS Sustainable and Green Finance Institute (SGFIN), NUS Asian Institute of Digital Finance (AIDF), Ministry of Education, Singapore under its MOE Academic Research Fund Tier 2 (STEM RIE2025 Award MOE-T2EP20123-0005: “Neurosymbolic AI for Commonsense-based Question Answering in Multiple Domains”), MOE Tier 2 Award (MOE-T2EP50221-0006: “Prediction-to-Mitigation with Digital Twins of the Earth’s Weather”), MOE Tier 1 Award (MOE-T2EP50221-0028: “Discipline-Informed Neural Networks for Interpretable Time-Series Discovery”), and by the RIE2025 Industry Alignment Fund – Industry Collaboration Projects (IAF-ICP) (Award I2301E0026: “Generative AI"), administered by A*STAR, as well as supported by Alibaba Group and NTU Singapore.